# Decompositional Reasoning for Graph Retrieval with Large Language Models

### v1

@@ 6.2 Graph size
$K_{n}$ and $K_{e}$ correspond respectively to the number of relevant nodes and edges that we consider to build the connected subgraph with PCST. At the retrieval step, we set the values of $K_{n}$ and $K_{e}$ to extract a certain number of relevant entities in the original graph (Figure 3). Choosing higher values of $K_{n}$ and $K_{e}$ leads to a higher quantity of retrieved information, which improves the probability of retrieved relevant nodes and edges, but also increases potential noise in the subgraph that we are building. Logically, choosing higher values of $K_{n}$ produces significantly larger graphs (Figure 12 in Appendix A), which can be harder to handle for the LLM.
We also show (see Figure 7) that setting a high value for $K_{n}$ (or $K_{e}$) does not lead to better performances for our 7B model. This observation has also been made in He et al. (2024b) on the WebQSP dataset. Setting $K_{n}$ too low (e.g. $K_{n}=2$) does not allow the model to retrieve enough knowledge in the graph; but setting $K_{n}$ too high (above 5 empirically for our method) will add noise (non-relevant nodes) to the retrieved subgraphs and can disrupt the correctness of the generated answers. If we use a larger model (see Figure 8), the difference between using $K_{n}=3$, $K_{e}=5$) and $K_{n}=5$, $K_{e}=7$ ( denoted as "larger graphs") does not lead to significant improvement; although we have a higher change of retrieving important nodes and edges, this observation highlights the presence of noise within the larger retrieved graphs. For the evaluation of our method, we use the default values of $K_{n}=3$ and $K_{e}=5$.
@@ 6.3 Main Results
Ours (Hybrid 7B/13B) 57.9 50.3 77.9 58.2
Ours (LLaMa-2-13B) 58.1 50.8 77.4 56.4
@@
For our main evaluation, we consider various baselines and model configurations. In particular, we highlight the "Hybrid 7B/13B" setting, where a 7B model answers each subquestion and a 13B model handles final answer generation (as described in Section 5.3).
Across both CWQ and WebQSP benchmarks (Table 1), our method achieves strong performance compared to approaches using similar model sizes and no fine-tuning. On CWQ, which features multi-hop questions, we observe a significant improvement over prior non-finetuned baselines, including those using larger models like Sun et al. (2024) (70B) and He et al. (2024b) (13B). On WebQSP, a simpler QA dataset, our method still outperforms related methods, though the margin is smaller—likely because decomposition is less helpful for single-hop questions. In both cases, only methods relying on dataset-specific fine-tuning or very large models (e.g., GPT-3.5 in (Chen et al., 2024)) achieve better scores, highlighting the value of simple decompositional reasoning at the retrieval stage. A key observation is that our "Hybrid 7B/13B" setup performs comparably to a full 13B pipeline, suggesting that most of the benefits come from decompositional retrieval, not simply model scale. Figure 8 highlights this efficiency: we maintain competitive performance while using fewer resources, by relying on a lightweight model for subquestions and a larger one only for the final answer.
Finally, Table 2 compares the average number of LLM-calls for our method and compares it with baselines (Sun et al. (2024), Chen et al. (2024)) that made this data available. These methods use iterative cycles to answer the complex question, which does not give any upper-bound for the number of calls to the model. In our case, the number of calls to the model directly depends on the number of generated subquestions, which can ultimately be controlled via prompting at the decomposition step. We achieve state-of-the-art accuracy while reducing LLM usage for both CWQ and WebQSP, showing the efficiency of our decompositional retrieval method.
@@
Since we use a single LLM call for both decomposition and final answer generation, we can deduce the average number of subquestions generated. Without setting a limit on the number of subquestions, we obtained an average of 2.8 subquestions for CWQ and 2.3 for WebQSP; this demonstrates that more complex questions result in more subquestions.
@@ 7 Conclusion
In this work, we introduced a novel graph retrieval method using decompositional reasoning with LLMs. By leveraging textual knowledge graphs and a hybrid retrieval mechanism that balances subquestion-specific and global context, our method enhances both the accuracy and interpretability of multi-hop QA. We demonstrated the effectiveness of our approach on complex QA benchmarks such as CWQ and WebQSP, achieving improved performance over strong baselines without increasing model size. Our results highlight the value of structured knowledge and explicit reasoning steps in addressing the limitations of LLMs in knowledge-intensive tasks. Future work may explore adding reasoning mechanisms at the generation step for improved model capabilities.
@@
Although our method demonstrated state-of-the-art results with smaller LLMs, we can mention some limitations of our method. Our method is mostly adapted to complex QA datasets, as the decomposition works best on difficult and multi-hop questions that can be transformed into a set of simple and atomic questions. The decomposition is not systematic, and we prompt the LLM to not decompose a question if it is considered to be simple enough; this approach can work on simple QA datasets (shown in Table 1 for the WebQSP dataset), but there is no guaranty that the model will not force the decomposition of simple questions.
We decompose complex questions using an LLM with a specific prompting technique shown in Appendix B. This method is advantageous for preprocessing an entire dataset, but it requires the use of a large enough LLM (we use Deepseek-R1-Distill-Qwen-32B, which is still relatively small compared to other baselines used for direct reasoning). Also, it is hard to control the quality of the decomposition; manual evaluation has been conducted to control the quality of the decomposition. It has been observed that some generated subquestions were redundant or irrelevant to the final goal, which can act as noise when providing them to the model.
@@ Appendix A Experimental Results
The value of the $\alpha$ parameter, which controls the hybrid retrieval mechanism, can cause the retrieved graphs to be more or less connected. We see on Figure 9 that with a lower value of $\alpha$, we sometimes produce disconnected graphs; at a higher value of $\alpha$, most (if not all) graphs become naturally connected. Figure 5 suggests that the model better handles the connected graphs, as they lead to better results, but the low number of disconnected graphs questions the statistical significance of this hypothesis. We observe that for some alpha values, the p-value is less than 0.05. We also use a Beta law to estimate the posterior distribution of p, the parameter for the Binomial law that represents the accuracy of our predictions. Over the different alpha values, we obtain $P(connected>disconnected)=0.919$, which indicates the plausibility of our claim.

### v3

@@ 6.2 Graph size
According to Figure 3, we can control the size of the retrieved nodes (noted as $K_{n}$) and edges (noted as $K_{e}$) for each subgraph. At the retrieval step, we set the values of $K_{n}$ and $K_{e}$ to extract a certain number of relevant entities in the original graph (Figure 3). Choosing higher values of $K_{n}$ and $K_{e}$ leads to a higher quantity of retrieved information, which improves the probability of retrieving relevant nodes and edges, but also increases the noise in the subgraph that we are building. We show that higher values of $K_{n}$ and $K_{e}$ produces significantly larger graphs (Figure 11 in Appendix A), which is harder to handle for the LLM. We show (Appendix A for detailed findings) that setting $K_{n}$ and $K_{e}$ too low will not allow retrieving enough relevant information, while setting them to high introduces noise in the final subgraph. For larger models (Figure 7), using larger graphs ($K_{n}=5$, $K_{e}=7$) offers marginal gains over smaller graphs ($K_{n}=3$, $K_{e}=5$), further indicating noise in larger subgraphs. We use $K_{n}=3$ and $K_{e}=5$ as default values for evaluation.
@@ 6.3 Main Results
DAGR (Hybrid 7B/13B) 57.9 50.3 77.9 58.2
DAGR (LLaMa-2-13B) 58.1 50.8 77.4 56.4
@@
For our main evaluation, we consider various baselines and model configurations. In particular, we highlight our "Hybrid 7B/13B" setting, where a 7B model answers each subquestion and a 13B model handles final answer generation (as described in Section 5.3).
Across both CWQ and WebQSP benchmarks (Table 1), DAGR achieves strong performance compared to approaches using similar model sizes and no fine-tuning. On CWQ, which features multi-hop questions, we observe a significant improvement over prior non-finetuned baselines, including those using larger models like Sun et al. (2024) (70B) and He et al. (2024b) (13B). On WebQSP, a simpler QA dataset, our method still equals most related methods, although decomposition might be less helpful for single-hop questions. In both cases, only methods relying on dataset-specific fine-tuning or very large models (e.g., GPT-3.5 in (Chen et al., 2024)) achieve better scores, highlighting the value of simple decompositional reasoning at the retrieval stage. A key observation is that our "Hybrid 7B/13B" retrieval setup performs similarly to a full 13B pipeline, suggesting that most of the benefits come from decompositional retrieval, not simply model scale. Figure 7 highlights this efficiency: we maintain competitive performance while using fewer resources, by relying on a lightweight model for subquestions and a larger one only for the final answer.
@@
Finally, Table 2 compares the average number of LLM-calls for our method and compares it with baselines that made this data available (Sun et al. (2024), Chen et al. (2024)). These methods use iterative cycles to answer the complex question, which does not give any upper-bound for the number of calls to the model. In our case, the number of calls to the model directly depends on the number of generated subquestions, which can ultimately be controlled via prompting at the decomposition step. While achieving state-of-the-art accuracy, we also notably reduce LLM usage for both datasets, showing the efficiency of our decompositional retrieval method. Since we only use a single LLM call for both decomposition and final answer generation, we can deduce the average number of subquestions generated. Without setting a limit on the number of subquestions, we obtained an average of 2.8 subquestions for CWQ and 2.3 for WebQSP, which shows that more complex questions will be decomposed in more subquestions.
@@ 7 Conclusion
We propose DAGR, a graph retrieval method using decompositional reasoning with LLMs, combining textual knowledge graphs and hybrid retrieval to improve multi-hop QA. Our approach achieves state-of-the-art accuracy on standard benchmarks without increasing model size, highlighting the value of structured knowledge and explicit reasoning in knowledge-intensive tasks.
@@
Although our method demonstrated state-of-the-art results with smaller LLMs, we can mention some limitations of the approach. Our method is mostly adapted to complex QA datasets, as the decomposition works best on difficult and multi-hop questions that can be transformed into a set of simple and atomic questions. The decomposition is not systematic, and we prompt the LLM to not decompose a question if it is considered to be simple enough; this approach can work on simple QA datasets (shown in Table 1 for the WebQSP dataset), but there is no guaranty that the model will not force the decomposition of simple questions.
We decompose complex questions using an LLM with a specific prompting technique shown in Appendix B. This method is advantageous for preprocessing an entire dataset, but it requires the use of a large enough LLM (we use Deepseek-R1-Distill-Qwen-32B, which is still relatively small compared to other baselines used for direct reasoning). Also, it is hard to control the quality of the decomposition; manual evaluation has been conducted to control the quality of the decomposition. It has been observed that some generated subquestions were redundant or irrelevant to the final goal, which can act as noise when providing them to the model. Potential future work might focus on implementing a error detection mechanism to control the quality of generated subquestions ; a path to explore would be to use a "LLM-as-a-judge" technique, where a larger model could be used to judge the quality of the generated subquestions.
@@ Appendix A Experimental Results
The value of the $\alpha$ parameter, which controls the hybrid retrieval mechanism, can cause the retrieved graphs to be more or less connected. We saw (Figure 5) that with a lower value of $\alpha$, we sometimes produce disconnected graphs; at a higher value of $\alpha$, most (if not all) graphs become naturally connected. Figure 8 suggests that the model better handles the connected graphs, as they lead to better results, but the low number of disconnected graphs questions the statistical significance of this hypothesis. We observe that for some alpha values, the p-value is less than 0.05. We also use a Beta law to estimate the posterior distribution of p, the parameter for the Binomial law that represents the accuracy of our predictions. Over the different alpha values, we obtain $P(connected>disconnected)=0.919$, which indicates the plausibility of our claim.
@@
Figure 12 highlights the importance of choosing the right values for $K_{n}$ at retrieval. Increasing $K_{n}$ or $K_{e}$ beyond certain values does not improve performance, which is consistent with findings from He et al. (2024b) on other datasets. Setting $K_{n}$ too low (e.g., $K_{n}=2$) limits knowledge retrieval, while higher values (e.g., $K_{n}>5$) introduce noise from irrelevant nodes, degrading answer quality.
@@
We test evaluate the quality of generated subquestions ; in that end, we perform a manual evaluation for a random sample of 200 complex questions. We classify each instance in one of four classes ; the results of the evaluation are presented in Table 4:
Quality 7B 14B 32B
++ (Excellent) 27 34 69.5
+ (Good) 11 11 12.5
- (Weak) 27 24.5 11.5
- - (Poor) 35 30.5 7.5
Table 4: Evaluation of question decomposition quality across four levels: Excellent (++), Good (+), Weak (–), and Poor (– –). Evaluation was perfomed a random sample of 200 questions. Results are expressed in percentages.
Our evaluation is made in the following way: an excellent decomposition (++) corresponds to a human-level decomposition ; a good decomposition (+) corresponds to a satisfying decomposition that a human subject would slightly change ; a weak decomposition (-) represents cases where some subquestions repeat or are not useful, which can cause disturbance, but will likely not break the reasoning chain ; finally, a poor decomposition (- -) corresponds to a false or non-existing decomposition. We observe that the "smaller" models produce a decomposition which is much less qualitative. There is a large gap between the 14B and 32B models, highlighting the emerging capabilities of LLMs with a larger number of parameters.
Decomposition Model 7B 14B 32B
Accuracy (Hit@1) 49.2 49.9 54.9
Table 5: We measure the accuracy (Hit@1) obtained using different model sizes at the decomposition step ; for the inference pipeline, we use LLaMa-2-7B. Results show the importance in the quality of the decomposition.
We can analyze how the noise in the subquestions propagates through the inference pipeline and the consequences on the overall performance: in Table 5, we observe that that the performance of the overall pipeline decreases when using a smaller model at decomposition. This statement confirms the manual evaluation that was done previously, where we observed a lower quality in the generated subquestions when using 7B and 14B models. In the manual evaluation, we saw that the 7B and 14B models produced subquestions of similar quality, while the 32B model outperformed them quite clearly ; in terms of accuracy, wa find similar results, as the 7B and 14B models lead to similar accuracy, and the 32B model outperforms them by more than 5%.